two identical incandescent light bulbs, each with resistance R=2Ω, are connected to as source with E= 8V and negligible internal resistance. find the current through each bulb is the bulb are connected in series.


1
Expert's answer
2021-05-04T06:45:39-0400

"R_T=R_1+R_2"

"R_T = 2+2 = 4 ohms"

For resistors in series current across the resistors is equal.

"I= \frac{V}{R}"

Where

V- Voltage

I- Current

R- Resistance

"I = \frac {8}{4}= 2A"

Current across each resistor is equal to 2A